OpenStudy (anonymous):

If I have the circular equation (x+2)squared + (y-7)squared = 196, what is the coordinates for the center?

OpenStudy (jdoe0001):

hmm

OpenStudy (jdoe0001):

\(\bf (x-{\color{brown}{ h}})^2+(y-{\color{blue}{ k}})^2=r^2\qquad center\ ({\color{brown}{ h}},{\color{blue}{ k}})\)

OpenStudy (anonymous):

\[\bf (x-{\color{brown}{ h}})^2+(y-{\color{blue}{ k}})^2=r^2\qquad center\ ({\color{brown}{ h}},{\color{blue}{ k}})\] \[\bf (x-{\color{brown}{ (-2)}})^2+(y-{\color{blue}{ 7}})^2=r^2\qquad center\ ({\color{brown}{ red~number}},{\color{blue}{ blue~number}})\]

OpenStudy (anonymous):

@jdoe0001 invisible friend gave out the formula without any explanation :) I supported him to make it understandable. :-)